OpenStudy (anonymous):

Help please! A toy rocket was launched from the ground. The function f(x) = -16x2 + 256x shows the height of the rocket f(x), in feet, from the ground at time x seconds. What is the axis of symmetry of the graph of f(x) and what does it represent? (3 points)

OpenStudy (anonymous):

x = 8; it takes 8 seconds to reach the maximum height and 8 seconds to fall back to the ground x = 16; it takes 16 seconds to reach the maximum height and 16 seconds to fall back to the ground x = 8; it takes 8 seconds to reach the maximum height and 16 seconds to fall back to the ground x = 16, it takes 16 seconds to reach the maximum height and 32 seconds to fall back to the ground

OpenStudy (anonymous):

@mathway

OpenStudy (anonymous):

Alright well we see this is an upside down parabola... lets find the vertex (where it changes direction) −b2a=−1282(−16)=12832=4 so the 'x' coordinate of our vertex is 4... so I'm going to say the rocket would take 4 seconds to get up to maximum height...and then it will fall back down taking 4 seconds How do i know it will only take 4 seconds to get down? Well I would find out the x-intercepts of the parabola by factoring -16x^2 + 128x = 0 factor our like terms -16x(x - 8) = 0 so we have 2 equations -16x = 0 and (x - 8) = 0 well that means our 2 x intercepts are 0 and 8 So if it take 4 seconds to go up to the max height...we only have 4 seconds to fall back down Hence the answer being B

OpenStudy (anonymous):

thank you!!!

OpenStudy (anonymous):

your welcome! :)

OpenStudy (anonymous):

bc i wasnt sure how to find it

OpenStudy (anonymous):

my sister had the same question yesterday so her answer was b but im not sure which answer choice is 4 for you lol :D

OpenStudy (anonymous):

okay i may need help again late

OpenStudy (anonymous):

ok feel free to tag me :)